Use algebra of sets to prove that,

 [(𝐵 − 𝐴)' ∩ 𝐴] − 𝐴' = 𝐴


1
Expert's answer
2021-10-19T08:02:58-0400

Solution:

"LHS=[(𝐵 − 𝐴)' ∩ 𝐴] − 𝐴'
\\=[(𝐵' − 𝐴')∩ 𝐴] − 𝐴'"

"=[(𝐵' − 𝐴') ∩ 𝐴] ∩ [𝐴']' \ \ [\because P-Q=P∩Q']"

"=[(𝐵' − 𝐴') ∩ 𝐴] ∩ A
\\=(𝐵' − 𝐴') ∩ 𝐴∩ 𝐴
\\=(𝐵' − 𝐴') ∩ 𝐴"

"\\=(𝐵 − 𝐴)' ∩ 𝐴
\\=A-(B-A)
\\=A
\\=RHS"

Hence, proved.